OpenStudy (anonymous):

BASIC MATHEMATICS AND ELEMENTARY ALGEBRA: A house worth 6.25M was given a discount rate of 20%. What additional rate of discount must be given to reduce the price to 4M?

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow [6,250,000(1 - 0.20)] - 4,000,000 }\)

OpenStudy (anonymous):

6.25*80/100*x/100=4 solve for x and it will give you 80 ..subtract it from 100 it will give 20..so additional 20% discountis needed

Parth (parthkohli):

See, the thing within square brackets is the formula for discount, and then after you find it, then you have to subtract that to find the additional price.

Parth (parthkohli):

Am I right?

OpenStudy (anonymous):

\[6.25 \times (80/100) \times ((100 - x)/100) = 4 \implies x = 20\]

OpenStudy (anonymous):

i dont get the answer?

OpenStudy (anonymous):

are you looking for the answer only?

OpenStudy (anonymous):

no. i want to learn too, but i just dont get it